# Two-stage Design for Failure Probability Estimation with Gaussian Process Surrogates

### v2

@@ 3.2 Hybrid Monte Carlo for failure probability estimation
Now for $\mathbf{x}_{i}\in\mathcal{U}$, upgrade surrogate prediction $\mu_{n}(\mathbf{x}_{i})$ to expensive evaluation of the true $f(\mathbf{x}_{i})$, resulting in the following “hybrid MC” estimator:
@@
The hybrid nature of this estimator is in the use of both true simulator evaluations (left term) and surrogate evaluations (right term). The accuracy of this estimator is contingent upon the classification accuracy of the surrogate and the choice of $\mathcal{U}$. As long as we restrict the size of $\mathcal{U}$ so only $B$-many samples from $X_{M}$ fall within it, i.e., $\sum_{i=1}^{M}\mathbbm{1}_{\left\{\mathbf{x}_{i}\in\mathcal{U}\right\}}=B$, then we will meet our budgetary constraints.
Rather than thinking of $\mathcal{U}$ as a continuum, we find it easier to conceptualize in the context of the discrete set $X_{M}$. In a perfect world, $\mathcal{U}$ would capture all the samples which the surrogate would misclassify. We are not privy to this information, and we are under a limited budget, but we can leverage the surrogate’s UQ to identify the $B$-many samples at which it is most uncertain. Recall, classification entropy (5) quantifies the degree of uncertainty in the pass/fail prediction of a surrogate. Consequently, our proposed allocation of samples to $\mathcal{U}$ is strikingly simple – place the $B$-many samples from $X_{M}$ with the highest entropy into $\mathcal{U}$. Denote these as $X_{M}^{B}$. Revisiting Figure 2, notice the high entropy regions (green dotted/dashed) align well with the misclassification regions (yellow shading). The entropy criterion is effectively identifying regions of high classification uncertainty where we could most utilize expensive true simulator evaluations to correct inaccurate surrogate predictions.
Ultimately, after evaluating $f(\mathbf{x})$ for $\mathbf{x}\in X_{M}^{B}$, we update the surrogate with these additional observations, thereby updating $\mu_{n}(\mathbf{x})$, before returning $\hat{\alpha}_{\textrm{HYB}}$ (7). Since the $X_{M}^{B}$ points are likely near a failure contour, they could significantly improve the accuracy of the surrogate predictions for other samples near the contour. For reference, consider the right panel of Figure 3. The black circles showing the contour locating design for the GP surrogate are unchanged, but here the additional budget of 50 evaluations is spent on the MC samples from $p(\mathbf{x})$ with the highest classification entropy (blue triangles). Compared to the left panel which continues to explore the entire domain, these highest entropy samples exclusively exploit the predicted failure contour. Compared to the bias samples portrayed in the middle panel, where the original surrogate design is abandoned, our two-stage design is able to leverage all simulator evaluations in its final estimate.
@@ 3.3 Summary and related work
We have presented a comprehensive strategy for estimating Eq. (1) from a budget of size $N$. It is summarized as follows. Start with a very large Monte Carlo sample from $p(\mathbf{x})$, denoted $X_{M}$. Train a surrogate model through a contour locating sequential design, monitoring the progress of the $\hat{\alpha}_{\textrm{SURR}}$ estimate for given $X_{M}$, and halting contour location once two consecutive updates are within ±$\pm\hat{\sigma}_{\alpha}$. Denote the design points chosen in this stage as $X_{n}$. Then, use the trained surrogate to identify the highest entropy samples, $X_{M}^{B}$, evaluate the simulator at these inputs, and update the surrogate with these new observations. Finally, return $\hat{\alpha}_{\textrm{HYB}}$ following Eq. (7).
@@
Given the deterministic nature of $f(\mathbf{x})$, effective surrogates should interpolate observed points. In this light, the hybrid MC estimator (7) is akin to a surrogate MC estimator (4) when the surrogate has been trained on all $\mathbf{x}_{i}\in\mathcal{U}$. This connection permits us to frame our contribution as a two-stage design. In the first stage, $X_{n}$ is chosen over $\mathcal{X}$ based on contour location. In the second stage, $X_{M}^{B}$ is chosen over $\mathbbm{P}_{\mathbf{x}}$ based on maximum entropy. The final design of size $N$ is then $X_{N}=[X_{n}^{\top},{X_{M}^{B}}^{\top}]^{\top}$.
@@
There are some noteworthy distinctions between these two design stages. To illustrate, refer back to the black circles ($X_{n}$) and blue triangles ($X_{M}^{B}$) in the right panel of Figure 3. In the first stage, the $n$ points selected for contour location may be acquired anywhere in the domain $\mathcal{X}$, seeking to balance both exploration of the response surface and exploitation of the suspected contour. In fact, the contour locating sequential design strategies we consider in Section 4 (Cole et al.,, 2023; Booth et al.,, 2024b) intentionally circumvent acquisitions based on maximal entropy in order to avoid clustered acquisitions and to promote more exploration of the surface. Alternatively, the $B$ points chosen in the second stage must originate from $p(\mathbf{x})$ and are solely exploitative. By nature, they will be clustered around the predicted contour. We contend that having a mix of data from these two realms is superior to either extreme.
@@
Some of the works we referenced in Section 1 have investigated stopping criteria for surrogate training in reliability settings. The most similar to our proposed criterion is that of Zhu and Du, (2016) who recommend a convergence criterion akin to $\frac{\hat{\sigma}_{\alpha}}{\hat{\alpha}}\leq\epsilon$, where $\epsilon$ is a user specified threshold. While similar to our stopping criterion, the choice of $\epsilon$ here is make-or-break, and little guidance is provided. If $\epsilon$ is too high, contour location will be halted too soon; if too low, contour location will continue past the point of utility. Our proposed stopping criterion avoids this choice and may be easily automated without problem-specific tuning. In our synthetic experiments, we find that it consistently results in appropriate termination with budget to spare.
@@
Hybrid MC estimators (similar to Eq. 7) have been entertained in other settings. Li and Xiu, (2010) use polynomial chaos models for a hybrid MC estimator with samples delegated to $\mathcal{U}$ based on predicted proximity to the contour, i.e., $|\mu_{n}(\mathbf{x}_{i})-t|$. This work, however, does not consider surrogate design and does not leverage surrogate UQ. Fuhrländer and Schöps, (2020) use hybrid MC, but they only consider training the surrogate on samples from $p(\mathbf{x})$. Echard et al., (2011) similarly train only on a subset of $X_{M}$. We attempted to recreate the approach of Fuhrländer and Schöps, (2020) in our setting: (i) sample large $X_{M}$, (ii) train an initial surrogate on a random selection of these samples, (iii) use the surrogate with a metric such as entropy to select additional batches from $X_{M}$ for evaluation through the simulator, (iv) update the surrogate, and (v) repeat steps (iii)-(iv) until the budget is exhausted. Yet we were unable to obtain even reasonable results because initial surrogate fits were very poor, and surrogate-informed acquisitions were unable to recover. Random selections from $X_{M}$ were overwhelmingly in high probability regions, not low probability failure regions. Hristov et al., (2019) noticed this conundrum and stated it well: “using a surrogate that is built on frequent events to predict and analyze rare events” can be futile.
@@
To visualize this conundrum in a simple setting, we revisit the Herbie example of Figure 1. With the input distribution centered at the origin, the probability of generating any samples outside $[-1,1]^{2}$ is very low. But a GP surrogate would benefit from some “anchor points” in these outer regions (Gramacy and Apley,, 2015). This distinction will be less prominent if $p(\mathbf{x})$ is uniformly distributed over $\mathcal{X}$, but the trade-off between exploring in the training phase and exploiting in the estimation phase will still be impactful. Even with uniform $p(\mathbf{x})$, training a surrogate only on maximal entropy points is suboptimal (Cole et al.,, 2023; Booth et al.,, 2024b).
@@ 4 Synthetic experiments
In this section we benchmark our method against state-of-the-art competitors on a variety of synthetic experiments. We consider four functions: the 2d Herbie function (Lee et al.,, 2011) with a total budget of 150, the 3d Ishigami function (Ishigami and Homma,, 1990) with a total budget of 300, the 6d Hartmann function (Picheny et al.,, 2013) with a total budget of 600, and the 4d Plateau function (Booth et al.,, 2024b) with a total budget of 200. The first three of these are plausibly stationary, so we use a standard GP surrogate, fit with the Scikit-learn package in python (Pedregosa et al.,, 2011), and the entropy contour location scheme of Cole et al., (2023). The Plateau function is known to be nonstationary; it is characterized by flat regions with a steep sloping drop between them. To address the nonstationarity, we use a Bayesian DGP surrogate (Sauer et al.,, 2023b), fit with the deepgp package in R (Booth,, 2023), with the contour location scheme of Booth et al., (2024b). With the right surrogate, the failure contour of the Plateau function can be well estimated from a rather small amount of data.
@@
Plateau $4.308\times 10^{-4}$ 4 30 200 $3.5\times 10^{6}$ DGP Booth et al., (2024b)
@@
First, we deploy our stopping criterion (Section 3.1) for each of the four scenarios, starting with an initial LHS sample of size $n_{0}$ (Table 1) and repeating for 30 re-randomized Monte Carlo repetitions. To avoid premature termination, we require a minimum of 10 observed failures and $n\geq 2n_{0}$ (atleast as many contour locating acquisitions as original space-filling observations). Figure 5 shows the distribution of chosen $n$ values. Notice, all trials terminated before the allotted $N$, reserving some evaluations for additional estimation, but there was considerable variability across trials. We view this variability as a strength – some random initializations need more data to effectively locate the contour while some get lucky from the start. Allowing the surrogate to dictate its own budget mitigates some of the risks of having an unlucky starting design.
For each trained surrogate with chosen $n$ (which varies across repetitions), we collect a large Monte Carlo sample ($X_{M}$, randomized across repetitions), then deploy the following estimation methods:
  • All CL: Ignoring the chosen $n$, continue contour location until the entire budget is exhausted and return the surrogate MC estimate (4).
  • Hybrid Entropy: Our two-stage approach using the hybrid MC estimate (7) where $\mathcal{U}$ is chosen to contain the $B$ points from $X_{M}$ with highest entropy.
  • Hybrid Proximity: Same as “Hybrid Entropy,” but $\mathcal{U}$ is chosen to contain the $B$ points from $X_{M}$ which are predicted to be closest to the failure threshold (i.e., the samples with the smallest $|\mu_{n}(\mathbf{x}_{i})-t|$). This is inspired by the work of Li and Xiu, (2010).
  • SIIS: Surrogate-informed importance sampling (6) with a GMM bias distribution (Peherstorfer et al.,, 2016). The bias distribution is trained on the predicted failures, i.e., $\left\{\mathbf{x}_{i}\in X_{M}\;|\;\mu_{n}(\mathbf{x}_{i})>t\right\}$. For our GP surrogates implemented in python, we use the scikit-learn package (Pedregosa et al.,, 2011) to train GMM bias distributions. For our DGP surrogates implemented in R, we use the mclust package (Scrucca et al.,, 2016).
@@
Each method uses the same $X_{M}$ and the same surrogate trained up to point $n$. The distinction is in what is done with the remaining $B$ samples. As implemented, the latter three are to some degree an internal comparison since they use our proposed stopping criterion to determine the budget allocation between surrogate training and estimation (Figure 5). Without borrowing from our proposed approach, there would be no guidance on the allocation of $n$ and $B$ for these methods. The “Hybrid Proximity” approach is truly an internal comparison as it borrows everything from our proposed two-stage design aside from the specific allocation of samples to $\mathcal{U}$. It is included to show the impact of the entropy selection criterion, which uses surrogate UQ, over a selection solely based on the predicted mean. Reproducible code for all experiments, including both the python and R implementations, is available in our public git repository. [^2] https://bitbucket.org/boothlab/failprob/
@@
Across the board, both importance sampling based methods (“SIIS” and “SIIS UCB”) suffered from high variability in their failure probability estimates. Although their median performance was generally on target (indicative of their unbiasedness), these methods consistently offered the poorest performance. The more conservative “SIIS UCB” had no apparant advantage over the traditional “SIIS.” We believe importance sampling is simply ill-suited for our low data settings. In all scenarios aside from the simple two-dimensional Herbie function, our two-stage design (“Two-stage Entropy”) outperformed the exhaustive contour location approach (“All CL”). Particularly as dimension grows and the response surface becomes more complex, there is a clear advantage to reserving some simulator evaluations for the second stage. Most of the time there was agreement between the two criteria for allocating samples to $\mathcal{U}$ (“Entropy” and “Proximity”), but when there was disagreement the entropy criterion offered higher accuracy, with uniformly superior worst-case performance. As suspected, the entropy criterion which incorporates uncertainty is more effective at identifying samples which the surrogate would misclassify.

### v3

@@ 3.2 Stage 2: Hybrid Monte Carlo for failure probability estimation
Now for $\mathbf{x}_{i}\in\mathcal{U}$, we upgrade surrogate prediction $\mu_{n}(\mathbf{x}_{i})$ to expensive evaluation of the true $f(\mathbf{x}_{i})$, resulting in the following “hybrid MC” estimator:
@@
The hybrid nature of this estimator is in the use of both true simulator evaluations (left term) and surrogate evaluations (right term). The accuracy of this estimator is contingent upon the classification accuracy of the surrogate and the choice of $\mathcal{U}$. If we restrict the size of $\mathcal{U}$ so only $B$-many samples from $X_{M}$ fall within it, i.e., $\sum_{i=1}^{M}\mathbbm{1}_{\left\{\mathbf{x}_{i}\in\mathcal{U}\right\}}=B$, then we can use this hybrid estimator while staying within our budgetary constraints. The evaluation of the expensive simulator at the Monte Carlo samples within $\mathcal{U}$ constitutes the second stage of our design.
Rather than thinking of $\mathcal{U}$ as a continuum, we find it easier to conceptualize in the context of the discrete set $X_{M}$. In a perfect world, $\mathcal{U}$ would capture all the samples which the surrogate would misclassify. We are not privy to this information, and we are under a limited budget, but we can leverage the surrogate’s UQ to identify the $B$-many samples at which it is most uncertain. Recall, classification entropy (5) quantifies the degree of uncertainty in the pass/fail prediction of a surrogate. Consequently, our proposed allocation of samples to $\mathcal{U}$ is strikingly simple – place the $B$-many samples from $X_{M}$ with the highest entropy into $\mathcal{U}$. Denote these as $X_{M}^{B}$. Revisiting Figure 3, notice the high entropy regions (green) align well with the misclassification regions (yellow shading). The entropy criterion is effectively identifying regions of high classification uncertainty where we could most utilize expensive true simulator evaluations to correct inaccurate surrogate classifications.
Ultimately, after completing the second design stage by evaluating $f(\mathbf{x})$ for $\mathbf{x}\in X_{M}^{B}$, we update the surrogate with these additional observations, thereby updating $\mu_{n}(\mathbf{x})$, before returning $\hat{\alpha}_{\textrm{HYB}}$ (7). The process of observing the true simulator at the high entropy locations and updating the surrogate accordingly could be done in small batches within a feedback loop as demonstrated in Figure 1. Yet in our exercises, we have not seen smaller batch sizes to improve performance (we provide empirical evidence of this in Supplementary Material). Accordingly, we prefer to collect a single batch of all $B$-many high entropy samples at once.
@@ 3.3 Summary and related work
We have presented a comprehensive strategy for estimating Eq. (1) from a budget of size $N$. It is summarized as follows. Start with a very large Monte Carlo sample from $p(\mathbf{x})$, denoted $X_{M}$. Train a surrogate model through a contour locating sequential design, monitoring the progress of the $\hat{\alpha}_{\textrm{SURR}}$ estimate (4) for given $X_{M}$, and halting contour location once two consecutive updates are within ±$\pm\hat{\sigma}_{\alpha}$. Denote the design points chosen in this stage as $X_{n}$. Then, use the trained surrogate to identify the highest entropy samples, $X_{M}^{B}$, evaluate the simulator at these inputs, and update the surrogate with these new observations. Finally, return $\hat{\alpha}_{\textrm{HYB}}$ following Eq. (7).
@@
Given the deterministic nature of $f(\mathbf{x})$, effective surrogates should interpolate observed points, i.e., $\mu_{n}(\mathbf{x}_{i})=f(\mathbf{x}_{i})$ when the surrogate has been trained on $\{\mathbf{x}_{i},y_{i}\}$. In this light, the hybrid MC estimator (7) is akin to a surrogate MC estimator (4) when the surrogate has been trained on all $\mathbf{x}_{i}\in\mathcal{U}$. This connection permits us to frame our contribution as a two-stage design. In the first stage, $X_{n}$ is chosen over $\mathcal{X}$ based on contour location. In the second stage, $X_{M}^{B}$ is chosen over $\mathbbm{P}_{\mathbf{x}}$ based on maximum entropy. The final design of size $N$ is then $X_{N}=[X_{n}^{\top},{X_{M}^{B}}^{\top}]^{\top}$.
@@
There are some noteworthy distinctions between these two design stages, as depicted in Figure 1. In the first stage, the $n$ points selected for contour location may be acquired anywhere in the domain $\mathcal{X}$, seeking to balance both exploration of the response surface and exploitation of the suspected contour. In fact, the contour locating sequential design strategies we consider in Section 4 (Cole et al.,, 2023; Booth et al.,, 2024c) intentionally circumvent acquisitions based on maximal entropy in order to avoid clustered acquisitions and to promote more exploration of the surface. Alternatively, the $B$ points chosen in the second stage must originate from $p(\mathbf{x})$ and are solely exploitative. By nature, they will be clustered around the predicted contour, where pass/fail classifications are hardest to pin down. We contend that having a mix of data from these two realms is superior to either extreme.
@@
Some of the previous works referenced in Section 1 have investigated stopping criteria for surrogate training in reliability settings. The most similar to our proposed criterion is that of Zhu and Du, (2016) who recommend a convergence criterion akin to $\frac{\hat{\sigma}_{\alpha}}{\hat{\alpha}}\leq\epsilon$, where $\epsilon$ is a user-specified threshold. While similar to our stopping criterion, the choice of $\epsilon$ here is make-or-break, and little guidance is provided. If $\epsilon$ is too high, contour location will be halted too soon; if too low, contour location will continue past the point of utility. Our proposed stopping criterion avoids this choice and may be easily automated without problem-specific tuning. In our synthetic experiments, we find that it consistently results in appropriate termination with budget to spare.
@@
Hybrid MC estimators (similar to Eq. 7) have been entertained in other settings. Li and Xiu, (2010) use polynomial chaos models for a hybrid MC estimator with samples allocated to $\mathcal{U}$ based on predicted proximity to the contour, i.e., $|\mu_{n}(\mathbf{x}_{i})-t|$. Their work, however, does not consider surrogate design and does not leverage surrogate UQ. Fuhrländer and Schöps, (2020) use hybrid MC, but they only consider training the surrogate on samples from $p(\mathbf{x})$. Echard et al., (2011) similarly train only on a subset of $X_{M}$. We attempted to recreate the approach of Fuhrländer and Schöps, (2020) in our setting: (i) sample large $X_{M}$, (ii) train an initial surrogate on a random selection of these samples, (iii) use the surrogate with a metric such as entropy to select additional batches from $X_{M}$ for evaluation through the simulator, (iv) update the surrogate, and (v) repeat steps (iii)-(iv) until the budget is exhausted. Yet we were unable to obtain even reasonable results because initial surrogate fits were very poor, and surrogate-informed acquisitions were unable to recover. Random selections from $X_{M}$ were overwhelmingly in high probability regions, not low probability failure regions. Hristov et al., (2019) noticed this conundrum and stated it well: “using a surrogate that is built on frequent events to predict and analyze rare events” can be futile.
@@
To visualize this conundrum in a simple setting, we revisit the Herbie example of Figure 2. With the input distribution centered at the origin, the probability of generating any samples outside $[-1,1]^{2}$ is very low. But a GP surrogate would benefit from some “anchor points” in these outer regions (Gramacy and Apley,, 2015). This distinction will be less prominent if $p(\mathbf{x})$ is uniformly distributed over $\mathcal{X}$, but the trade-off between exploring in Stage 1 and exploiting in Stage 2 will still be impactful. Even with uniform $p(\mathbf{x})$, training a surrogate only on maximal entropy points is suboptimal (Cole et al.,, 2023; Booth et al.,, 2024c).
@@ 4 Synthetic experiments
In this section we benchmark our method against state-of-the-art competitors on a variety of synthetic experiments. We consider four functions: the 2d Herbie function (Lee et al.,, 2011) with a total budget of 150, the 3d Ishigami function (Ishigami and Homma,, 1990) with a total budget of 300, the 6d Hartmann function (Picheny et al.,, 2013) with a total budget of 600, and the 4d Plateau function (Booth et al.,, 2024c) with a total budget of 200. The first three of these are plausibly stationary, so we use a standard GP surrogate, fit with the Scikit-learn package in python (Pedregosa et al.,, 2011), and the entropy contour location scheme of Cole et al., (2023). The Plateau function is known to be nonstationary; it is characterized by flat regions with a steep sloping drop between them. To address the nonstationarity, we use a Bayesian DGP surrogate (Sauer et al.,, 2023b), fit with the deepgp package in R (Booth,, 2024), and the contour location scheme of Booth et al., (2024c). With the right surrogate, the failure contour of the Plateau function can be well estimated from a rather small amount of data.
@@
Plateau $4.308\times 10^{-4}$ 4 30 200 $3.5\times 10^{6}$ DGP Booth et al., (2024c)
@@
First, we deploy our stopping criterion (Section 3.1) for each of the four scenarios (Table 1). Starting with an initial LHS sample of size $n_{0}$, we proceed with Stage 1, conducting CL with the specified method and determining the $n$ at which to stop. We conduct 30 re-randomized Monte Carlo repetitions. To avoid premature termination, we require a minimum of 10 observed failures and $n\geq 2n_{0}$ (at least as many contour locating acquisitions as original space-filling observations). Figure 5 shows the distribution of chosen $n$ values for each scenario. Notice, all trials terminated before the allotted $N$, reserving some evaluations for Stage 2, but there was considerable variability across trials. We view this variability as a strength – some random initializations need more data to effectively locate the contour while some get lucky from the start. Allowing the surrogate to dictate its own budget allocation mitigates some of the risks of having an unlucky starting design.
For each trained surrogate with chosen $n$ (which varies across repetitions), we collect a large Monte Carlo sample ($X_{M}$, constant across methods, randomized across repetitions), then compare the following estimation methods. All variations use the same Stage 1 contour location scheme up to point $n$; they differ only in their use of the remaining $B=N-n$ observations.
  • Exhaustive CL: Ignoring the chosen $n$, continue contour location until the entire budget is exhausted and return the surrogate MC estimate (4).
  • Two-stage Entropy: Our proposed two-stage approach (Figure 1) using the hybrid MC estimate (7) where $\mathcal{U}$ is chosen to contain the $B$ points from $X_{M}$ with highest entropy.
  • Two-stage Proximity: Same as “Two-stage Entropy,” but $\mathcal{U}$ is chosen to contain the $B$ points from $X_{M}$ which are predicted to be closest to the failure threshold (i.e., the samples with the smallest $|\mu_{n}(\mathbf{x}_{i})-t|$). This is inspired by the work of Li and Xiu, (2010).
  • SIIS: Surrogate-informed importance sampling (6), with $B$-many samples observed from an estimated GMM bias distribution (Peherstorfer et al.,, 2016). The bias distribution is trained on the predicted failures, i.e., $\left\{\mathbf{x}_{i}\in X_{M}\;|\;\mu_{n}(\mathbf{x}_{i})>t\right\}$. For our GP surrogates implemented in python, we use the scikit-learn package (Pedregosa et al.,, 2011) to train GMM bias distributions. For our DGP surrogates implemented in R, we use the mclust package (Scrucca et al.,, 2016).
@@
As implemented, the latter three are to some degree an internal comparison since they use our proposed stopping criterion to determine the budget allocation between surrogate training ($n$) and estimation ($B$). Without borrowing from our proposed approach, there would be no guidance on the allocation of $n$ and $B$ for these methods. The “Two-stage Proximity” approach is truly an internal comparison as it borrows everything from our proposed two-stage design aside from the specific allocation of samples to $\mathcal{U}$. It is included to show the impact of the entropy selection criterion, which uses surrogate UQ, over a selection solely based on the posterior mean. Reproducible code for all experiments, including both the python and R implementations, is available in our public git repository. [^2] https://bitbucket.org/boothlab/failprob/
@@
Across the board, both importance sampling based methods (“SIIS” and “SIIS UCB”) suffered from high variability in their failure probability estimates. Although their median performance was generally on target (indicative of their unbiasedness), these methods consistently offered the poorest performance. The more conservative “SIIS UCB” had no apparent advantage over the traditional “SIIS.” We believe importance sampling is simply ill-suited for our low data settings. In all scenarios aside from the simplest two-dimensional Herbie function (where performance was comparable), our two-stage design (“Two-stage Entropy”) outperformed the exhaustive contour location approach (“Exhaustive CL”). Particularly as dimension grows and the response surface becomes more complex, there is a clear advantage to our two-stage design which transitions from CL to exploitation of MC samples. Within the two-stage designs, there was generally agreement between the two variations we considered for allocating samples to $\mathcal{U}$ (“Entropy” and “Proximity”), but when there was disagreement the entropy criterion offered higher accuracy, with uniformly superior worst-case performance. As suspected, the entropy criterion which incorporates uncertainty is more effective at identifying samples which the surrogate would misclassify.